OpenStudy (anonymous):

Please help! Show that the surfaces z=1/2 (x^2 + y^2 - 1) and z=1/2(1 - x^2 - y^2) are orthoganal at all points of intersection (We know two surfaces are said to be orthogonal to each other at a point P if the normals to their tangent planes are perpendicular at P. )

OpenStudy (anonymous):

Take the gradient of each curve. That is the normal vector of their respective tangent planes.

OpenStudy (anonymous):

Then take the dot product of the two gradients and show that you'll get 0.

OpenStudy (anonymous):

Ok.. so how do I go about finding the normal vector of the tangent plane?

OpenStudy (anonymous):

Have you taken up gradients?

OpenStudy (anonymous):

No we haven't

OpenStudy (anonymous):

How about partial derivatives?

OpenStudy (anonymous):

yeah!

OpenStudy (anonymous):

Ok the equation for the tangent plane for z=f(x,y) is z = f(x0,y0) + fx (x-x0) + fy (y-y0) where (x0,y0) is a point on the surface. We can infer that the normal vector N = <fx, fy, -1> so use that to find the normal vectors and show that they are orthogonal

OpenStudy (anonymous):

Ok wait, what does that have to do with partial derivitives?!

OpenStudy (anonymous):

oh wait, nvm. so, would i just randomly pick a point on the surface .. or leave it as f(x0,y0) ?

OpenStudy (anonymous):

Pick a point. Actually you can compute vectorN=<fx, fy, -1> right away.

OpenStudy (anonymous):

what do you mean right away?

OpenStudy (anonymous):

Like what I said before, we can infer that the normal vector N=<fx,fy,-1>, no need to find a point and go through computing the equation of the plane.

OpenStudy (anonymous):

that's the normal for the first equation right?

OpenStudy (anonymous):

And the second is perpendicular to that? N = <1, -fy, -fx> ?

OpenStudy (anonymous):

For any equation. Let the first surface be z=f(x,y), then vector N1=<fx,fy,-1> Let the second be z=g(x,y), then vector N2=<gx,gy,-1>

OpenStudy (anonymous):

Ok, so z = f(x0, y0) + 2x (x-x0) + 2y (y-y0) and then z= f(x0,y0) -2x (x-x0) - 2y (y-y0)

OpenStudy (anonymous):

Use g instead of f for the 2nd surface to avoid confusion. Remember the normal vector of a plane ax+by+cz=d is just <a,b,c> or ai+bj+ck?

OpenStudy (anonymous):

Ok wait, I'm kind of confused now. So where do I go from here? z = f(x0, y0) + 2x (x-x0) + 2y (y-y0) z= g(x0,y0) -2x (x-x0) - 2y (y-y0)

OpenStudy (anonymous):

It should be 1st plane: z = f(x0,y0) + x(x-x0) + y(y-y0) 2nd plane: z = g(x0,y0) - x(x-x0) - y(y-y0) there is a (1/2) so fx=x, fy=y, gx=-x, gy=-y. Rearrange it to the form ax+by+z=constant 1st plane: x (x-x0) + y (y-y0) - z = -f(x0,y0) 2nd plane: -x (x-x0) - y (y-y0) - z = -g(x0,y0) So the components of the normal vectors are the coefficients of x,y,z. N1 = xi + yj -k N2 = -xi -yj -k (Do you use the ijk notation?)

OpenStudy (anonymous):

Take the dot/inner product. N1 dot N2 = -x^2 -y^2 +1 --> show this is =0

OpenStudy (anonymous):

so -x^2 - y^2 + 1 dot x^2 + y^2 - 1 = 0 ?

OpenStudy (anonymous):

The problem says at all points of intersection. so f(x,y,)=g(x,y) to get all the ponts (x,y) where the surfaces intersect. (1/2)(x^2+y^2-1)=(1/2)(-x^2-y^2+1) blah... blah... solve... solve.. we get x^2+y^2=1

OpenStudy (anonymous):

now N1 dot N2 = -x^2 -y^2 +1 = -(x^2 + y^2) +1 = -(1) + 1 = 0, therefore orthogonal at every point (x,y) of intersection.

OpenStudy (anonymous):

Oh ok! Thanks!

OpenStudy (anonymous):

dot product of the normal vectors.

OpenStudy (anonymous):

what? dot product of the normal vectors? thats just what you were saying before right?

OpenStudy (anonymous):

yeah. coz u said "so -x^2 - y^2 + 1 dot x^2 + y^2 - 1 = 0 ?" It should be (xi+yj-k) dot (-xi-yj-k)